Bustech’s standard terms on sales invoices are 30 days, but analysis has shown that on average customers pay on 60 days.

Total sales revenue is $9,000,000 and the company expects irrecoverable receivables of $90,000.

A suggestion has been made to offer a 10% discount to customers paying when ordering (ie 0 days). It is expected that 50% of customers will take advantage of this. Irrecoverable receivables would be expected to fall by $45,000 and administrative savings on credit control would amount to $35,000.

Bustech finances its working capital with a bank overdraft that costs 10% per annum.

a) Calculate the total savings and the total costs, and recommend whether or not the proposal should be considered.

b) Calculate the maximum discount that could be offered to be viable.
Answer #1

Answer to (a)

Analysis of data indicates that proposal should be considered as net outflow in proposal is less than the current setup. Calculations are as below:

Particulars Option 1 Option 2 Calculation (Option 1) Calculation (Option 2) Remarks
Total costs
Bad debts          90,000      45,000 As given As given
Interest on working capital          13,315         6,658 =(900000-90000)*10%*60/365 =((900000*50%)-45000)*10%*60/365 (Credit sales - Bad debts) X Interest rate
Discount expense                   -        45,000 =900000*50%*10% Cash sales X Discount rate
Total savings
Admin savings on credit control                   -      (35,000) As given
Net costs       103,315      61,658

Answer to (b)

Cost saving in Option 2 ignoring discount:

Particulars Option 1 Option 2 Calculation (Option 1) Calculation (Option 2) Remarks
Total costs
Bad debts          90,000      45,000 As given As given
Interest on working capital          13,315         6,658 =(900000-90000)*10%*60/365 =((900000*50%)-45000)*10%*60/365 (Credit sales - Bad debts) X Interest rate
Total savings
Admin savings on credit control                   -      (35,000) As given
Net costs       103,315      16,658

Savings = 103315 - 16658 = 86658

Therefore, maximum discount that can be given for option to be viable is $86658

Discount percentage = Discount amount / Cash sales = 86658/ (900000 X 50%)

= 19.26%